OpenStudy (electronia):

George has a square sandbox in the backyard. the sandbox has an area of 160 square feet. to the nearest foot, what is the length of one side of the sandbox? 11, 12,40,80

OpenStudy (steve816):

Hey, I hope you are having a wonderful day today. Remember that the area of a square is length times width, but since the lengths are equal, you can simply do\[\Large160 =x^2\] \[\Large x=\sqrt{160}\]

OpenStudy (electronia):

Oh so the square of 160 is the answer?

OpenStudy (steve816):

Yup!

OpenStudy (electronia):

So 12 (rounded)

OpenStudy (steve816):

You are correct :)

OpenStudy (electronia):

Thank you!

OpenStudy (steve816):

No problem!